# Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands

### v1

@@ Abstract
Non-analytic Bloch eigenstates at isolated band degeneracy points exhibit singular behavior in the quantum metric. Here, a description of superfluid weight for zero-energy flat bands in proximity to other high-energy bands is presented, where they together form a singular band gap system. When the singular band gap closes, the geometric and conventional contributions to the superfluid weight as a function of the superconducting gap exhibit different crossover behaviors. The scaling behavior of superfluid weight with the band gap is studied in detail, and the effect on the Berezinskii-Kosterlitz-Thouless (BKT) transition temperature is explored. It is discovered that tuning the singular band gap provides a unique mechanism for enhancing the supercurrent and critical temperature of two-dimensional (2D) superconductors.
@@
Introduction.—Flat bands or regions in momentum space with a large density of states [1, 2, 3, 4, 5], like saddle point Van Hove singularities [6, 7] might be promising for the realization of high-$T_{c}$ superconductors. However, due to the vanishing band curvature in 2D flat bands, phase fluctuations can drive the critical temperature $T_{c}$ to near zero, even for strong pairing interactions [8]. One mechanism for this is the BKT transition, which occurs due to vortex proliferation at temperatures $T>T_{c}$, resulting in resistive motions destroying superfluidity. In this case $T_{c}$ is determined from the Nelson-Kosterlitz criteria: $T_{c}=(\pi/8)D_{s}(T_{c}^{-})$ where $D_{s}$ denotes the superfluid weight [9, 10]. Therefore, it is critical to identify band structure properties that can enhance the superfluid weight and establish conditions for achieving maximum values of $T_{c}$.
@@
With the advent of moire superconductors [11, 12, 13, 14, 15], the role of band geometry in the superconducting fluctuations has been studied intensively [5, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 21, 28, 29]. In particular, it has been established that the superfluid weight of an isolated flat band is proportional to the integral of quantum metric, which is lower bounded by topological invariants. Therefore, band geometric properties determine the BKT transition temperature $T_{c}$ of flat bands [5, 16]. However, this analysis requires the flat band to be separated from other bands by a band gap $E_{g}$. In this Letter, we study the effect of the band gap $E_{g}$ on the superfluid weight of flat bands in proximity to other bands. We explicitly show its intriguing crossover behavior as a function of the superconducting gap $\Delta$ when $E_{g}$ is varied in a class of flat bands called singular flat bands [30, 31, 32]. This behavior is due to the non-analyticity of Bloch functions in singular flat bands, which can be used to enhance $T_{c}$ of flat-band superconductors.
Figure 1: Illustration of the crossover of geometric superfluid weight $D_{s}^{\text{geo}}$ as a function of superconducting gap $\Delta$, in some fixed energy unit, as the band gap $E_{g}$ between a singular flat and dispersive band changes. In the weak interaction case (indicated by red arrow), as $E_{g}$ closes $D_{s}^{\text{geo}}$ crossovers from $\propto\Delta$ to $\Delta\ln(W/2\Delta)$, with $W\approx 2$ the width of the dispersive band. The strong interaction case (black arrow) crossovers from $\propto\Delta$ to $1/\Delta$. The two dashed curves represent the $\propto\Delta$ and $1/\Delta$ asymptotic behaviors for the blue curve $E_{g}=1.5$. Inset: the conventional superfluid weight $D_{s}^{\text{conv}}$ vs $\Delta$. In the weak interaction regime, it crossovers from $\propto\Delta^{2}$ to $\Delta$. All the crossovers above occur at the regime $E_{g}\sim\Delta$.
@@
Singular flat bands exhibit essential (immovable) singularities in the Bloch eigenstates [31] resulting from the destructive interference of Bloch phases due to band touching degeneracies [30, 32]. Here, we introduce two classes of two- and three-band continuum Hamiltonians that host singular flat bands, separated from other high-energy bands of width $W_{d}$ by a tunable band gap $E_{g}$. Both classes of models have a momentum-space vortex with nonzero winding numbers at the gap closing point and are continuum model extensions of the Mielke [33] and Lieb [34] lattices. For these lattice systems, it was found that $D_{s}$ scales as $U\ln(\epsilon_{c}/U)$ due to the $k^{-2}$ divergence of quantum metric near the band touching point [16, 35, 36, 37], where $U>0$ is the attractive interaction and $\epsilon_{c}$ is a constant. We show this logarithmic enhancement of $D_{s}$ is a general feature of singular flat bands. In the weak interaction limit, defined by $\Delta\ll W_{d}$, the geometric contribution to the superfluid weight, $D_{s}^{\text{geo}}$ exhibits a crossover from $\Delta$ to $\Delta\ln(W/2\Delta)$ ($W$ is an energy scale of the order of $W_{d}$) as $E_{g}$ closes, with the crossover at $E_{g}\sim\Delta$, as indicated by the red arrow in Fig 1. In contrast, the conventional contribution $D_{s}^{\text{conv}}$ has a $\Delta^{2}$ to $\Delta$ crossover, as shown in the inset of Fig 1. In the strong interaction limit $\Delta\gg W_{d}$, $D_{s}^{\text{conv}}$ is suppressed, while $D_{s}^{\text{geo}}$ has the crossover from $\Delta$ to $1/\Delta$, indicated by the black arrow in Fig 1. The different scaling limits of the weak and strong interactions result in a maximum superfluid weight, which can be used to enhance the BKT transition temperature.
@@
where $\zeta_{J}$ is a constant with $J$ the integer chirality index, $m$ is the mass, and $k$, $\varphi_{\textbf{k}}$ are the magnitude and azimuth angle of the 2D momentum. Pauli matrices $\hat{\sigma}_{i}$ act on the orbital degree of freedom of the model. We impose time-reversal symmetry and assume zero spin-orbit coupling, which dictates that $h(\textbf{k})$ has a spin-down time-reversal partner. The spectrum of Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands includes a flat valence band $\varepsilon_{v\textbf{k}}=0$ and a dispersive conduction band $\varepsilon_{c\textbf{k}}=2(\zeta_{J}^{2}k^{2J}+m^{2})^{1/2}$. The bands acquire a winding number ±$\pm J$ upon opening the band gap $E_{g}=2m$ and are topological.
Figure 2: Momentum-space vortex of the vector field $\hat{\textbf{h}}({\bf k})$ for model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands of $J=1$. (a) $m=0$, $\hat{\textbf{h}}({\bf k})$ is frustrated at the origin. (b) $m\neq 0$, $\hat{\textbf{h}}({\bf k})$ gets tilted out of the $k_{x}-k_{y}$-plane.
Non-analyticity of Bloch eigenstates near the gap-closing point can be visualized by plotting the unit vector field $\hat{{\bf h}}(\textbf{k})={\bf h}(\textbf{k})/|{\bf h}(\textbf{k})|$, with $\textbf{h}(\textbf{k})=(h_{x},h_{y},h_{z})$ the vector associated with matrices $\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}$ in Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands. It exhibits a momentum-space vortex due to the winding number $J$, as indicated in Fig. 2. When $m=0$ (Fig. 2 a), $\hat{{\bf h}}(\textbf{k})$ is frustrated at the origin due to the non-analyticity of the Bloch states. However, when $m\neq 0$ (Fig. 2 b), $\hat{{\bf h}}(\textbf{k})$ cants out of the k-plane at the origin, thus is well-defined. The non-analyticity is also reflected in the quantum metric of both bands [38]:
@@
where $\hat{\lambda}_{i}$’s form a three-dimensional representation of the SU(2) algebra. The additional factor $C(m,k)$ is to allow for generic dispersive bands while preserving the chirality [39]. We focus on two limiting cases: (i) $C(m,k)=1$, the model is finite-hopping and consists of a zero-energy flat band (0) and two dispersive bands ($\pm$), with $\varepsilon_{\pm,\textbf{k}}=\pm(\zeta_{J}^{2}k^{2J}+m^{2})^{1/2}$; (ii) $C(m,k)=c(m)(\zeta_{J}^{2}k^{2J}+m^{2})^{-1/2}$, the dispersive bands are flattened by long-range hoppings and the function $c(m)$ depicts how the band gap depends on the mass $m$.
@@
For any choice of $C(m,k)$, model Eq. 3 has both particle-hole and inversion symmetry, implying a zero inter-band (band-resolved) quantum metric between the valence and conduction band, $g^{+-}_{\mu\nu}(\textbf{k})=0$, while the intra-band quantum metric $g^{0}_{\mu\nu},g^{\pm}_{\mu\nu}$ are all proportional to Eq. 2 (see supplemental materials). The middle flat band is topologically trivial, in contrast to the flat band of the two-band model. In case (i) when $C(m,k)=1$, the model of $J=1$ corresponds to the continuum limit of Lieb lattice, with a topological mass gap $m$ from a loop-ordered state [40]. It is distinct from the bipartite Lieb lattice model discussed in Ref. [16] where a band gap occurs due to unequal hoppings between the nearest neighbors. The relation of the two- and three-band models above to the singular flat bands previously studied in finite-hopping lattice models [30, 31, 41, 32] is provided in the supplemental materials.
@@
Weak interaction limit.—We first study the weak interaction limit ($\Delta\ll W_{d}$) of superfluidity for a system with flat bands at half-filling ($\mu\approx 0$) in proximity to other dispersive bands, as illustrated in Fig. 3. The system is characterized by three energy scales: the band gap $E_{g}$, the superconducting gap $\Delta$, and the cutoff energy $\varepsilon_{\Lambda}$ or width $W_{d}$ of the dispersive band. The effect of other energy scales, e.g., the width $W_{f}$ of the flat band with weak dispersion, is assumed to be subleading and neglected. The superconducting gap is characterized by a single parameter $\Delta$, consistent with the uniform pairing condition [5], i.e., all orbitals have the same pairing potential. We envision that the band gap $E_{g}$ can be tuned at a fixed electron density for the superconducting state. When $E_{g}\gg\Delta$ (Fig. 3(a)), the dispersive band has zero electron density. As $E_{g}$ approaches $\Delta$ (Fig. 3(b)) from above (below), the dispersive band becomes electron- (hole-) doped, shifting the chemical potential downwards (upward), $\delta\mu<0$ ($>0$). In the weak interaction regime $\Delta\ll\varepsilon_{\Lambda}$ or $W_{d}$, the change $\delta\mu$ is neglected as a first approximation. Since the density of states changes, the order parameter of each orbital $\Delta_{\alpha}$ ($\alpha$ is the orbital index) renormalizes with $E_{g}$. However, for chiral models, a moderate change of $E_{g}$ still corresponds to an approximate uniform pairing, allowing us to take the average of $\Delta_{\alpha}$ as $\Delta$ (see supplemental materials for details).
@@
where $\mathcal{F}_{1}^{\alpha},\mathcal{F}_{2}^{\alpha}$ are functions of dimensionless parameters $\lambda\equiv E_{g}/W$ and $\kappa\equiv W/\Delta$, with $W$ a phenomenological bandwidth of the dispersive band. Label $\alpha=``\text{conv}"$ or ``$\text{geo}"$ denotes the two contributions, and $\chi(\lambda,\kappa)$ is a crossover function, defined as
@@
$\lambda$ controls the band geometry of the gap system, whereas the quantity $\lambda\kappa=E_{g}/\Delta$ measures the magnitude of $E_{g}$ vs $\Delta$, determining the crossover.
We now explain the crossover from Eq. 4 above for the weak interaction regime $\Delta\ll W$ ($\kappa\gg 1$). When the band gap is large, $E_{g}\gg\Delta$ ($\lambda\kappa\gg 1$, Fig. 3(a)), $\kappa$ cancels in the logarithmic leading term of the crossover function so $\chi(\lambda,\kappa)\approx-\ln\lambda+...$, leading to the expansion of Eq. 4 in powers of $1/\kappa$:
@@
where $f_{1}^{\alpha},f_{2}^{\alpha}$ are the two leading coefficients. It shows that $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ scale as power laws of $\Delta$ when the flat band is isolated. Typically $f_{1}^{\text{geo}}\neq 0$ while $f_{1}^{\text{conv}}=0,f_{2}^{\text{conv}}\neq 0$ for perfectly flat bands, implying the behavior $D_{s}^{\text{geo}}\propto\Delta$ and $D_{s}^{\text{conv}}\propto\Delta^{2}$, respectively. Here, $D_{s}^{\text{conv}}$ is of higher order because the pairing at the edge of dispersive bands is extremely weak.
@@
This gives a $D_{s}^{\alpha}\sim\Delta\ln(W/2\Delta)$ behavior if $\mathcal{F}_{1}^{\alpha}(0,\kappa)\neq 0$, leading to a power law to logarithm crossover. However, if $\mathcal{F}_{1}^{\alpha}(0,\kappa)=0$, then Eq. 7 corresponds to another power law $D_{s}^{\alpha}\sim\Delta\mathcal{F}_{2}^{\alpha}(0,\kappa)$, and $\lambda\kappa\sim 1$ is the crossover regime between two power laws. We find $\mathcal{F}_{1}^{\text{conv}}(0,\kappa)=0$, leading to a $\Delta^{2}$ to $\Delta$ crossover for $D_{s}^{\text{conv}}$; whereas $\mathcal{F}_{1}^{\text{geo}}(0,\kappa)\neq 0$ due to the $k^{-2}$ divergence of quantum metric, leading to the distinctive $\Delta$ to $\Delta\ln(W/2\Delta)$ crossover for $D_{s}^{\text{geo}}$ in singular flat bands.
@@
Table 1: Functions $\mathcal{F}_{i}^{\alpha}$ and the leading coefficients $f_{i}^{\alpha}$ of the two-band chiral model Eq.Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands.
@@
Next, we address the physics of the two-band chiral model. Due to the combination of particle-hole and inversion symmetry, at half-filling, the superfluid weight of the three-band model in the dispersive case (i) exhibits the same functional form as the two-band model (see supplemental materials for details); therefore, the discussion below applies to it as well. Functions $\mathcal{F}^{\alpha}_{i}$, $f^{\alpha}_{i}$ of the two-band model are listed in Table 1. Here, the phenomenological bandwidth is defined as $W\equiv(\varepsilon_{\Lambda}^{2}+E_{g}^{2})^{1/2}$, with the cutoff energy $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$ and band gap $E_{g}=2m$. The chirality index $J$ is a multiplying factor, indicating that the crossover is universal for singular band-gap systems of different chiralities. The crossover can also be seen from the divergence of the leading coefficients $f_{1}^{\text{geo}}$ and $f_{2}^{\text{conv}}$ in Table 1 as $\lambda\rightarrow 0$—if they did not diverge, then no crossover would occur. Coefficients $f_{1}^{\text{geo}}$ and $f_{2}^{\text{conv}}$ diverge in the way of $\ln\lambda$ and $1/\lambda$, respectively, pointing to a nontrivial character of the geometric superfluid weight, which will be discussed shortly.
@@
Although we focused on the half-filling case, the crossover behavior persists at other fillings of flat bands as long as $\nu$ is not too close to 0 or 1. Away from half-filling, one must account for a nonzero $\mu$, which introduces another energy scale. Since the crossover occurs whenever the edge of the dispersive bands approaches the interaction window of $2\Delta$ around the chemical potential, it defines $|E_{g}-\mu|\sim\Delta$ as the crossover regime. However, the electron or hole density of flat bands must dominate for this. If $\nu$ is very close to 0 or 1, the chemical potential may lie in the dispersive bands, in which case the analysis above does not apply.

### v2

# Superfluid weight cross-over and critical temperature enhancement in singular flat bands

@@ Abstract
Non-analytic Bloch eigenstates at isolated band degeneracy points exhibit singular behavior in the quantum metric. Here, a description of superfluid weight for zero-energy flat bands in proximity to other high-energy bands is presented, where they together form a singular band gap system. When the singular band gap closes, the geometric and conventional contributions to the superfluid weight as a function of the superconducting gap exhibit different cross-over behaviors. The scaling behavior of superfluid weight with the band gap is studied in detail, and the effect on the Berezinskii-Kosterlitz-Thouless (BKT) transition temperature is explored. It is discovered that tuning the singular band gap provides a unique mechanism for enhancing the supercurrent and critical temperature of two-dimensional (2D) superconductors.
@@
Introduction.—Flat bands or regions in momentum space with a large density of states [1, 2, 3, 4, 5, 6], like saddle point Van Hove singularities [7, 8, 9, 10] might be promising for the realization of high-$T_{c}$ superconductors. However, due to the vanishing band curvature in 2D perfectly flat bands, phase fluctuations can drive the critical temperature $T_{c}$ to near zero, even for strong pairing interactions [11]. One mechanism for this is the BKT transition, which occurs due to vortex proliferation at temperatures $T>T_{c}$, resulting in resistive vortex motion destroying superfluidity. The BKT transition temperature $T_{c}$ of 2D superconductors is determined from the Nelson-Kosterlitz criteria: $T_{c}=(\pi/8)D_{s}(T_{c})$ where $D_{s}$ denotes the superfluid weight (superfluid stiffness) [12, 13], which defines the supercurrent $\textbf{J}_{s}$ as a response to the vector potential A as $\textbf{J}_{s}=-D_{s}\textbf{A}$. Therefore, it is critical to identify band structure properties that can enhance the superfluid weight and establish conditions for achieving maximum values of $T_{c}$.
@@
The superfluid weight in conventional BCS theory is inversely proportional to the effective mass, which vanishes for flat bands. This raises questions about the nature of superconductivity in moire superconductors hosting flat bands [14, 15, 16, 17, 18, 19]. In flat-band superconductors band geometric effects rescue the superconductivity, providing a finite superfluid weight [6, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 25, 32, 33, 34]. The superfluid weight of an isolated flat band is proportional to the integral of quantum metric, which is lower bounded by topological invariants. Therefore, band geometric properties determine the BKT transition temperature $T_{c}$ of flat-band superconductors [6, 20]. However, this analysis requires the flat band to be separated from other bands by a band gap $E_{g}$. In this Article, we study the effect of the band gap $E_{g}$ on the superfluid weight of flat bands in proximity to other bands. We show its intriguing cross-over behavior as a function of the superconducting gap $\Delta$ when $E_{g}$ is varied in a class of flat bands called singular flat bands [35, 36, 37]. This behavior is due to the non-analyticity of Bloch functions in singular flat bands, which can be used to enhance $T_{c}$ of flat-band superconductors.
@@
Singular flat bands exhibit essential (immovable) singularities in the Bloch eigenstates [36] resulting from the destructive interference of Bloch phases due to band touching degeneracies [35, 37]. Here, we introduce two classes of two- and three-band continuum Hamiltonians that host singular flat bands, separated from other high-energy bands of width $W_{d}$ by a tunable band gap $E_{g}$. Both classes of models have a momentum-space vortex with nonzero winding numbers at the gap closing point and are continuum model extensions of the Mielke [38] and Lieb [39] lattices. For these lattice systems, it was found that $D_{s}$ scales as $U\ln(\epsilon_{c}/U)$ due to the $k^{-2}$ divergence of the quantum metric near the band touching point [20, 40, 41, 42], where $U>0$ is the attractive interaction and $\epsilon_{c}$ is a constant. We show this logarithmic enhancement of $D_{s}$ is a general feature of singular flat bands. In the weak interaction limit, defined by $\Delta\ll W_{d}$, the geometric contribution to the superfluid weight, $D_{s}^{\text{geo}}$ exhibits a cross-over from $\Delta$ to $\Delta\ln(W/2\Delta)$ ($W$ is an energy scale of the order of $W_{d}$) as $E_{g}$ closes, with the cross-over at $E_{g}\sim\Delta$, as indicated by the red arrow in Fig 1. In contrast, the conventional contribution $D_{s}^{\text{conv}}$ has a $\Delta^{2}$ to $\Delta$ cross-over, as shown in the inset of Fig 1. In the strong interaction limit $\Delta\gg W_{d}$, $D_{s}^{\text{conv}}$ is suppressed, while $D_{s}^{\text{geo}}$ has the cross-over from $\Delta$ to $1/\Delta$, indicated by the black arrow in Fig 1. The different scaling limits of the weak and strong interactions result in a maximum superfluid weight, which can be used to enhance the BKT transition temperature.
Figure 1: Illustration of the cross-over of geometric superfluid weight $D_{s}^{\text{geo}}$ as a function of superconducting gap $\Delta$, in some fixed energy unit, as the band gap $E_{g}$ between a singular flat and dispersive band changes. In the weak interaction case (indicated by red arrow), as $E_{g}$ closes $D_{s}^{\text{geo}}$ cross-overs from $\propto\Delta$ to $\Delta\ln(W/2\Delta)$, with $W\approx 2$ the width of the dispersive band. The strong interaction case (black arrow) cross-overs from $\propto\Delta$ to $1/\Delta$. The two dashed curves represent the $\propto\Delta$ and $1/\Delta$ asymptotic behaviors for the blue curve $E_{g}=1.5$. Inset: the conventional superfluid weight $D_{s}^{\text{conv}}$ vs $\Delta$. In the weak interaction regime, it cross-overs from $\propto\Delta^{2}$ to $\Delta$. All the cross-overs above occur at the regime $E_{g}\sim\Delta$.
@@
where $\zeta_{J}$ is a constant with $J$ the integer chirality index, $m$ is the mass, and $k$, $\varphi_{\textbf{k}}$ are the magnitude and azimuth angle of the 2D momentum. Pauli matrices $\hat{\sigma}_{i}$ act on the orbital degree of freedom of the model. We impose time-reversal symmetry and assume zero spin-orbit coupling, which dictates that $h(\textbf{k})$ has a spin-down time-reversal partner. The spectrum of Eq. Superfluid weight cross-over and critical temperature enhancement in singular flat bands includes a flat valence band $\varepsilon_{v\textbf{k}}=0$ and a dispersive conduction band $\varepsilon_{c\textbf{k}}=2(\zeta_{J}^{2}k^{2J}+m^{2})^{1/2}$. The bands acquire a winding number ±$\pm J$ upon opening the band gap $E_{g}=2m$ and are topological.
Figure 2: Momentum-space vortex of the vector field $\hat{\textbf{h}}({\bf k})$ for model Eq. Superfluid weight cross-over and critical temperature enhancement in singular flat bands of $J=1$. (a) $m=0$, $\hat{\textbf{h}}({\bf k})$ is frustrated at the origin. (b) $m\neq 0$, $\hat{\textbf{h}}({\bf k})$ gets tilted out of the $k_{x}-k_{y}$-plane.
Non-analyticity of Bloch eigenstates near the gap closing point can be visualized by plotting the unit vector field $\hat{{\bf h}}(\textbf{k})={\bf h}(\textbf{k})/|{\bf h}(\textbf{k})|$, with $\textbf{h}(\textbf{k})=(h_{x},h_{y},h_{z})$ the vector associated with matrices $\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}$ in Eq. Superfluid weight cross-over and critical temperature enhancement in singular flat bands. It exhibits a momentum-space vortex due to the winding number $J$, as indicated in Fig. 2. When $m=0$ (Fig. 2 a), $\hat{{\bf h}}(\textbf{k})$ is frustrated at the origin due to the non-analyticity of the Bloch states. However, when $m\neq 0$ (Fig. 2 b), $\hat{{\bf h}}(\textbf{k})$ cants out of the k-plane at the origin, thus is well-defined. The non-analyticity is also reflected in the quantum metric of both bands [43]:
@@
where $\hat{\lambda}_{i}$’s form a three-dimensional representation of the SU(2) algebra. The additional factor $C(m,k)$ is to allow for generic dispersive bands while preserving the chirality  [^1] Considering the fixed electron density, we engineer the band structure by $C(m,k)$ for the three-band model only. For the two-band model, when both bands are flattened, the chemical potential $\mu=0$ (aligned with the valence band) and electron density cannot be fixed simultaneously, making the analytical analysis more difficult. For the three-band model, there is no such limitation.. We focus on two limiting cases: (i) $C(m,k)=1$, the model is finite-hopping and consists of a zero-energy flat band (0) and two dispersive bands ($\pm$), with $\varepsilon_{\pm,\textbf{k}}=\pm(\zeta_{J}^{2}k^{2J}+m^{2})^{1/2}$; (ii) $C(m,k)=c(m)(\zeta_{J}^{2}k^{2J}+m^{2})^{-1/2}$, the dispersive bands are flattened by long-range hoppings and the function $c(m)$ depicts how the band gap depends on the mass $m$.
@@
For any choice of $C(m,k)$, model Eq. 3 has both particle-hole and inversion symmetry, implying a zero inter-band (band-resolved) quantum metric between the valence and conduction band, $g^{+-}_{\mu\nu}(\textbf{k})=0$, while the intra-band quantum metric $g^{0}_{\mu\nu},g^{\pm}_{\mu\nu}$ are all proportional to Eq. 2 (see supplemental materials Sec. S3). The middle flat band is topologically trivial, in contrast to the flat band of the two-band model. In case (i) when $C(m,k)=1$, the model of $J=1$ corresponds to the continuum limit of Lieb lattice, with a topological mass gap $m$ from a loop-ordered state [45]. It is distinct from the bipartite Lieb lattice model discussed in Ref. [20] where a band gap occurs due to unequal hoppings between the nearest neighbors. The relation of the two- and three-band models above to the singular flat bands previously studied in finite-hopping lattice models [35, 36, 46, 37] is provided in the supplemental materials Sec. S1.
Superfluid Weight.—To analyze the energy cost of superconducting phase fluctuations in singular flat bands, we calculate the superfluid weight for onsite attractive interactions. For the models we study in this Article, the orbitals are spatially separated, therefore onsite attractive interactions lead to intra-orbital interactions of the form $\hat{V}=-\sum_{\alpha}U_{\alpha}\hat{n}_{i\alpha\uparrow}\hat{n}_{i\alpha\downarrow}$, where $i$ labels the unit cell, $\alpha$ labels the orbital, and $\hat{n}_{i\alpha\sigma}=c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$ is the density operator for spin $\sigma$. Then mean-field decoupling results in an intra-orbital order parameter: $\Delta_{\alpha}=-U_{\alpha}\langle c_{i\alpha\downarrow}c_{i\alpha\uparrow}\rangle$. With the uniform pairing condition $\Delta_{\alpha}\equiv\Delta$ for all orbitals $\alpha$ in the singular flat bands, the total superfluid weight $D_{s}=D_{s}^{\text{conv}}+D_{s}^{\text{geo}}$ for the two-band chiral model can be expressed as
$D_{s}=\sum_{{\bf k}}\frac{\Delta^{2}}{E_{c,{\bf k}}^{3}}(\partial_{x}\varepsilon_{c,{\bf k}})^{2}+4\Delta\sum_{{\bf k}}\bigg{(}1-\frac{\Delta}{E_{c,{\bf k}}}\bigg{)}g_{xx}({\bf k}),$ (4)
where $E_{c,{\bf k}}=(\varepsilon_{c,{\bf k}}^{2}+\Delta^{2})^{1/2}$ is the quasiparticle energy, and the sum of k is limited by the bandwidth $W_{d}$ (for derivations see supplementary materials Sec. S3). We have used the relation $D_{s,xx}=D_{s,yy}$ and $D_{s,xy}=0$ to define $D_{s}=D_{s,xx}$. Similar expressions of superfluid weight for the three-band chiral model with particle-hole and inversion symmetry can also be derived (see supplemental materials Sec. S3.1 for details). For some cases, we nominally adjusted the interactions $U_{\alpha}$ to ensure the uniform pairing condition.
@@
Weak interaction limit.—We first study the weak interaction limit ($\Delta\ll W_{d}$) of superfluidity for a system with flat bands at half-filling ($\mu\approx 0$) in proximity to other dispersive bands, as illustrated in Fig. 3. The system is characterized by three energy scales: the band gap $E_{g}$, the superconducting gap $\Delta$, and the cutoff energy $\varepsilon_{\Lambda}$ or width $W_{d}$ of the dispersive band. The effect of other energy scales, e.g., the width $W_{f}$ of the flat band with weak dispersion, is assumed to be subleading and neglected. The superconducting gap is characterized by a single parameter $\Delta$, consistent with the uniform pairing condition [6], i.e., all orbitals have the same pairing potential. We envision that the band gap $E_{g}$ can be tuned at a fixed electron density for the superconducting state. When $E_{g}\gg\Delta$ (Fig. 3(a)), the dispersive band has zero electron density. As $E_{g}$ approaches $\Delta$ (Fig. 3(b)) from above (below), the dispersive band becomes electron- (hole-) doped, shifting the chemical potential downwards (upward), $\delta\mu<0$ ($>0$). In the weak interaction regime $\Delta\ll\varepsilon_{\Lambda}$ or $W_{d}$, the change $\delta\mu$ is neglected as a first approximation. Since the density of states changes, the order parameter of each orbital $\Delta_{\alpha}$ ($\alpha$ is the orbital index) renormalizes with $E_{g}$. However, for chiral models, a moderate change of $E_{g}$ still corresponds to an approximate uniform pairing, allowing us to take the average of $\Delta_{\alpha}$ as $\Delta$ (see supplemental materials Sections S5, S6, S7 for details).
@@
where $\mathcal{F}_{1}^{\alpha},\mathcal{F}_{2}^{\alpha}$ are functions of dimensionless parameters $\lambda\equiv E_{g}/W$ and $\kappa\equiv W/\Delta$, with $W$ a phenomenological bandwidth of the dispersive band. Label $\alpha=``\text{conv}"$ or ``$\text{geo}"$ denotes the two contributions, and $\chi(\lambda,\kappa)$ is a cross-over function, defined as
@@
$\lambda$ controls the band geometry of the gap system, whereas the quantity $\lambda\kappa=E_{g}/\Delta$ measures the magnitude of $E_{g}$ vs $\Delta$, determining the cross-over.
We now explain the cross-over from Eq. 5 above for the weak interaction regime $\Delta\ll W$ ($\kappa\gg 1$). When the band gap is large, $E_{g}\gg\Delta$ ($\lambda\kappa\gg 1$, Fig. 3(a)), $\kappa$ cancels in the logarithmic leading term of the cross-over function so $\chi(\lambda,\kappa)\approx-\ln\lambda+...$, leading to the expansion of Eq. 5 in powers of $1/\kappa$:
@@
where $f_{1}^{\alpha},f_{2}^{\alpha}$ are the two leading coefficients. It shows that $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ scale as power laws of $\Delta$ when the flat band is isolated. Typically $f_{1}^{\text{geo}}\neq 0$ while $f_{1}^{\text{conv}}=0,f_{2}^{\text{conv}}\neq 0$ for perfectly flat bands, implying the behavior $D_{s}^{\text{geo}}\propto\Delta$ and $D_{s}^{\text{conv}}\propto\Delta^{2}$, respectively. Here $D_{s}^{\text{conv}}$ is of higher order because the pairing at the edge of the dispersive bands is extremely weak.
@@
This gives a $D_{s}^{\alpha}\propto\Delta\ln(W/2\Delta)$ behavior if $\mathcal{F}_{1}^{\alpha}(0,\kappa)\neq 0$, leading to a power law to logarithm cross-over. However, if $\mathcal{F}_{1}^{\alpha}(0,\kappa)=0$, then Eq. 8 corresponds to another power law $D_{s}^{\alpha}\propto\Delta\mathcal{F}_{2}^{\alpha}(0,\kappa)$, and $\lambda\kappa\sim 1$ is the cross-over regime between two power laws. We find $\mathcal{F}_{1}^{\text{conv}}(0,\kappa)=0$, leading to a $\Delta^{2}$ to $\Delta$ cross-over for $D_{s}^{\text{conv}}$; whereas $\mathcal{F}_{1}^{\text{geo}}(0,\kappa)\neq 0$ due to the $k^{-2}$ divergence of quantum metric, leading to the distinctive $\Delta$ to $\Delta\ln(W/2\Delta)$ cross-over for $D_{s}^{\text{geo}}$ in singular flat bands.
Table 1: Functions $\mathcal{F}_{i}^{\alpha}$ and the leading coefficients $f_{i}^{\alpha}$ of the two-band chiral model Eq.Superfluid weight cross-over and critical temperature enhancement in singular flat bands.
@@
Next, we address the physics of the two-band chiral model. Due to the combination of particle-hole and inversion symmetry, at half-filling, the superfluid weight of the three-band model in the dispersive case (i) exhibits the same functional form as the two-band model (see supplemental materials Sec. S3.1 for details); therefore, the discussion below applies to it as well. Functions $\mathcal{F}^{\alpha}_{i}$, $f^{\alpha}_{i}$ of the two-band model are listed in Table 1. Here, the phenomenological bandwidth is defined as $W\equiv(\varepsilon_{\Lambda}^{2}+E_{g}^{2})^{1/2}$, with the cutoff energy $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$ and band gap $E_{g}=2m$. The chirality index $J$ is a multiplying factor, indicating that the cross-over is universal for singular band-gap systems of different chiralities. The cross-over can also be seen from the divergence of the leading coefficients $f_{1}^{\text{geo}}$ and $f_{2}^{\text{conv}}$ in Table 1 as $\lambda\rightarrow 0$—if they did not diverge, then no cross-over would occur. Coefficients $f_{1}^{\text{geo}}$ and $f_{2}^{\text{conv}}$ diverge as $\ln\lambda$ and $1/\lambda$, respectively, pointing to a nontrivial character of the geometric superfluid weight, which will be discussed shortly.
@@
Although we focused on the half-filling case, the cross-over behavior persists at other fillings of flat bands as long as $\nu$ is not too close to 0 or 1. Away from half-filling, one must account for a nonzero $\mu$, which introduces another energy scale. Since the cross-over occurs whenever the edge of the dispersive bands approaches the interaction window of $2\Delta$ around the chemical potential, it defines $|E_{g}-\mu|\sim\Delta$ as the cross-over regime. However, flat bands’ electron or hole density must dominate for this. If $\nu$ is very close to 0 or 1, the chemical potential may lie in the dispersive bands, in which case the analysis above does not apply.